# ConsistCompose: Unified Multimodal Layout Control for Image Composition

### v1

@@ A.2 Stage 2: Hybrid Supervised Fine-Tuning
We further train ConsistCompose on a mixture of generation and understanding tasks, following the hybrid multi-task supervised fine-tuning (SFT) recipe used in Bagel [7]. On the generation side, the mixture comprises text-to-image, multi-reference-to-image, LELG-conditioned text-to-image, LELG-conditioned multi-reference-to-image, and image editing, with layout-conditioned tasks upsampled to reinforce layout consistency. On the understanding side, image-to-text pairs and interleaved understanding data are sampled at lower ratios so as to maintain general visual–language competence while keeping the overall training focused on layout-controllable generation.
@@ Appendix B Details of Coordinate-CFG
To improve layout controllability, we introduce a Coordinate-CFG mechanism that augments the text–image classifier-free guidance (CFG) used in Bagel [7] with an additional coordinate-guidance branch. When coordinate tokens are absent, the formulation reduces exactly to Bagel’s original two-branch CFG. Quantitative results on COCO-Position are reported in Table S6.
@@ B.2 Hierarchical fusion
This formulation is backward-compatible with the CFG used in Bagel [7]. When the coordinate-guidance branch is disabled at sampling time (i.e., $\mathbf{v}_{t}^{\text{final}}$ is computed from the text and image branches without the intermediate coordinate-fusion step), the update rule reduces to the original two-branch text–image CFG parameterized by $s_{\text{text}}$ and $s_{\text{img}}$. Disabling the image-guidance branch as well further recovers the standard text-only CFG.
@@ B.3 Recommended Coordinate-CFG range
We provide empirically validated coordinate-guidance scales based on MS-Bench [44] and COCO-Position evaluations.
@@ Appendix C Dataset Construction
(1) LayoutSAM. LayoutSAM [56] provides multi-object scenes with instance-level bounding boxes and textual descriptions. We convert its annotations into our unified training format and, for the layout-grounded text-to-image split, additionally supply a global scene description for each image that summarizes all entities and their relations.
(2) UNO-1M. UNO-1M [50] is built on top of Objects365 [34] and derives subject-centric data from its category set. We follow the class taxonomy used in UNO-1M [50]and directly reuse its processed outputs—subject reference images and attribute-aware phrases—as paired sets $\{R_{i}\}$ and $\{s_{i}\}$. No additional subject categories are introduced.
(3) Subjects200K. Subjects200K [36] is incorporated in the same manner, expanding both the coverage and appearance diversity of the subject bank.
(4) Virtual try-on data. To further enrich the reference-conditioned branch, we incorporate human–garment virtual try-on datasets such as VITON-HD  [5] and DressCode-MR [6]. Each sample consists of a reference garment image and a corresponding try-on image in which the garment is embedded within a layered human-body configuration, naturally inducing nested and occluded spatial relationships. We treat the garment image as the reference, annotate an axis-aligned bounding box for the garment region in the try-on image, and describe the scene using a standardized English virtual-try-on template.these samples provide complementary reference-conditioned supervision in which compositional constraints arise from body–garment interactions rather than explicit multi-object layouts.
@@ C.1 Two data construction branches
This branch is built from two sources: (i) LayoutSAM-style corpora [56] with given images, a global dense caption, and instance-level spatial annotations, and (ii) canonicalized samples derived from our reference-conditioned pipelines. For (i), LayoutSAM [56] provides, for each image, a global dense caption $p^{\text{raw}}_{\text{dense}}$ and a set of entities with class labels and bounding boxes, where each instance is already mentioned in $p^{\text{raw}}_{\text{dense}}$. We convert the global dense caption $p^{\text{raw}}_{\text{dense}}$ to a layout-grounded LELG prompt $p_{\text{LELG}}$, preserving its narrative content while explicitly encoding all annotated instances and their spatial coordinates, using GPT-4o in an in-context learning configuration with the prompt design shown in Fig. S10. For each instance, we append its normalized bounding-box tag $x_{1},y_{1},x_{2},y_{2}]</bbox$> immediately after the first occurrence of the corresponding phrase in $p_{\text{LELG}}$, producing layout-grounded T2I training pairs while transforming dense spatial annotations into inline, language-embedded supervision. For (ii), we start from the outputs of Pipelines 1–3 (described below) and apply the deterministic conversion in Sec. C.1.2: we drop all reference-image inputs and remove any “from image kk” indices from the tags, while keeping the scene sentence otherwise unchanged.
@@ C.1.2 Reference-conditioned branch
This branch is built from UNO-1M [50] and Subjects200K [36] subject banks and phrases processed by the three pipelines below (see Fig. S7). Each sample consists of a scene sentence with inline <bbox> tags and a set of subject reference images $\{R_{i}\}$. The pipelines are explicitly designed for subject-consistent, layout-guided generation. Through the canonicalization in Sec. C.1.1, the same samples also contribute to the layout-grounded T2I branch. In addition, we use virtual try-on data from VITON-HD [5] and DressCode-MR[6] as auxiliary sources, converting them into the same reference-conditioned format to further enrich the subject and appearance diversity of ConsistCompose3M.
Pipeline 1: Layout Diversification via Controlled Outpainting. As shown in Pipeline 1 of Fig. S7, this pipeline broadens the spatial distribution of bounding-box positions and scales in the single-reference setting using image pairs from Subjects200K [36], obtained through OmniControl’s side-by-side paired generation. For each sample, we first generate a dense caption $p^{\text{raw}}_{\text{dense}}$ for the reference image using GPT-4o [14]. The reference image $R_{i}$ is then resized by a sampled side-length factor $r\sim\mathcal{U}(0.6,0.8)$ and placed onto a blank canvas at a random valid location, thereby defining a normalized bounding box $b\in[0,1]^{4}$.
@@
Finally, the three pipelines collectively yield, for each constructed sample, a scene-level dense description and the corresponding subject–bounding-box associations. We then convert the raw dense sentence $p^{\text{raw}}_{\text{dense}}$ into a layout-grounded, multi-reference prompt $p_{\text{LELG}}$ by inserting explicit bounding-box and source-image tags. For each matched subject, we take its normalized bounding box $x_{1},y_{1},x_{2},y_{2}]$ together with the index of its reference image, and employ GPT-4o [14] with the instruction template in Fig. S12 to insert a tag of the form $x_{1},y_{1},x_{2},y_{2}]</bbox$> from image $k$ immediately after the first noun-phrase mention of that subject in $p^{\text{raw}}_{\text{dense}}$. By integrating reference-guided synthesis, detection-and-matching–based filtering, and LLM-driven tag insertion, these pipelines produce layout-grounded samples exhibiting rich nested and attributive configurations.
@@ Appendix D Supplementary Results and Detailed Analysis
General Understanding Capabilities. An important consideration is whether layout-aware training compromises general multimodal understanding, which we investigate through additional quantitative ablation studies across two benchmark suites. For fine-grained generation control (Table S7), our Alignment and SFT (incorporating LLM-based rewriting) attain an overall score of 0.88—ranking first among compared methods—with state-of-the-art position control performance (0.84) and strong attribute consistency (0.79–0.80). This confirms that precise layout adherence does not sacrifice control granularity. For general understanding benchmarks (Table S8), our Alignment variants (with/without coordinate data) achieve  81.4 on MMBench, comparable to InternVL3-8B (81.7) [63] and Bagel (81.4), validating no degradation in basic visual understanding capabilities. Notably, on MMMU (assessing complex reasoning), Alignment with coordinate data improves from 39.4 (without coordinate data) to 42.3. This suggests that coordinate data may enhance spatial reasoning, representing an advantage over specialized layout models that often trade generalizability for layout control.
@@
Extended DreamBench Results. To assess generalization beyond layout control, we additionally report DreamBench results (Tab. S9, Figs. S15 and S16). Alignment (w/ Coord Data) outperforms its non-coordinate counterpart (0.611 vs. 0.565 in the DreamBench single-subject DINO identity metric), and SFT (w/ Coord Data) further reaches 0.677 for single-subject and 0.506 for multi-subject generation, surpassing baselines such as OmniGen2 [49] and narrowing the gap to top-performing models (Flux Kontext dev [17], GPT-4o [14]). Qualitative examples in Figs. S15 and S16 corroborate these gains, showing strong identity preservation and prompt adherence for single subjects, as well as consistent identities with natural interactions in multi-subject scenes. Together, these results indicate that our method attains strong layout control without compromising general understanding or broad generative capability.

### v2

@@ A.2 Stage 2: Hybrid Supervised Fine-Tuning
We further train ConsistCompose on a mixture of generation and understanding tasks, following the hybrid multi-task supervised fine-tuning (SFT) recipe used in Bagel [deng2025emerging]. On the generation side, the mixture comprises text-to-image, multi-reference-to-image, LELG-conditioned text-to-image, LELG-conditioned multi-reference-to-image, and image editing, with layout-conditioned tasks upsampled to reinforce layout consistency. On the understanding side, image-to-text pairs and interleaved understanding data are sampled at lower ratios so as to maintain general visual–language competence while keeping the overall training focused on layout-controllable generation.
@@ Appendix B Details of Coordinate-CFG
To improve layout controllability, we introduce a Coordinate-CFG mechanism that augments the text–image classifier-free guidance (CFG) used in Bagel [deng2025emerging] with an additional coordinate-guidance branch. When coordinate tokens are absent, the formulation reduces exactly to Bagel’s original two-branch CFG. Quantitative results on COCO-Position are reported in Table S6.
@@ B.2 Hierarchical fusion
This formulation is backward-compatible with the CFG used in Bagel [deng2025emerging]. When the coordinate-guidance branch is disabled at sampling time (i.e., $\mathbf{v}_{t}^{\text{final}}$ is computed from the text and image branches without the intermediate coordinate-fusion step), the update rule reduces to the original two-branch text–image CFG parameterized by $s_{\text{text}}$ and $s_{\text{img}}$. Disabling the image-guidance branch as well further recovers the standard text-only CFG.
@@ B.3 Recommended Coordinate-CFG range
We provide empirically validated coordinate-guidance scales based on MS-Bench [wang2025msdiffusion] and COCO-Position evaluations.
@@ Appendix C Dataset Construction
(1) LayoutSAM. LayoutSAM [zhang2025creatilayout] provides multi-object scenes with instance-level bounding boxes and textual descriptions. We convert its annotations into our unified training format and, for the layout-grounded text-to-image split, additionally supply a global scene description for each image that summarizes all entities and their relations.
(2) UNO-1M. UNO-1M [wu2025less] is built on top of Objects365 [shao2019objects365] and derives subject-centric data from its category set. We follow the class taxonomy used in UNO-1M [wu2025less]and directly reuse its processed outputs—subject reference images and attribute-aware phrases—as paired sets $\{R_{i}\}$ and $\{s_{i}\}$. No additional subject categories are introduced.
(3) Subjects200K. Subjects200K [tan2025ominicontrol] is incorporated in the same manner, expanding both the coverage and appearance diversity of the subject bank.
(4) Virtual try-on data. To further enrich the reference-conditioned branch, we incorporate human–garment virtual try-on datasets such as VITON-HD  [choi2021viton] and DressCode-MR [chong2025fastfit]. Each sample consists of a reference garment image and a corresponding try-on image in which the garment is embedded within a layered human-body configuration, naturally inducing nested and occluded spatial relationships. We treat the garment image as the reference, annotate an axis-aligned bounding box for the garment region in the try-on image, and describe the scene using a standardized English virtual-try-on template.these samples provide complementary reference-conditioned supervision in which compositional constraints arise from body–garment interactions rather than explicit multi-object layouts.
@@ C.1 Two data construction branches
This branch is built from two sources: (i) LayoutSAM-style corpora [zhang2025creatilayout] with given images, a global dense caption, and instance-level spatial annotations, and (ii) canonicalized samples derived from our reference-conditioned pipelines. For (i), LayoutSAM [zhang2025creatilayout] provides, for each image, a global dense caption $p^{\text{raw}}_{\text{dense}}$ and a set of entities with class labels and bounding boxes, where each instance is already mentioned in $p^{\text{raw}}_{\text{dense}}$. We convert the global dense caption $p^{\text{raw}}_{\text{dense}}$ to a layout-grounded LELG prompt $p_{\text{LELG}}$, preserving its narrative content while explicitly encoding all annotated instances and their spatial coordinates, using GPT-4o in an in-context learning configuration with the prompt design shown in Fig. S10. For each instance, we append its normalized bounding-box tag $x_{1},y_{1},x_{2},y_{2}]</bbox$> immediately after the first occurrence of the corresponding phrase in $p_{\text{LELG}}$, producing layout-grounded T2I training pairs while transforming dense spatial annotations into inline, language-embedded supervision. For (ii), we start from the outputs of Pipelines 1–3 (described below) and apply the deterministic conversion in Sec. C.1.2: we drop all reference-image inputs and remove any “from image kk” indices from the tags, while keeping the scene sentence otherwise unchanged.
@@ C.1.2 Reference-conditioned branch
This branch is built from UNO-1M [wu2025less] and Subjects200K [tan2025ominicontrol] subject banks and phrases processed by the three pipelines below (see Fig. S7). Each sample consists of a scene sentence with inline <bbox> tags and a set of subject reference images $\{R_{i}\}$. The pipelines are explicitly designed for subject-consistent, layout-guided generation. Through the canonicalization in Sec. C.1.1, the same samples also contribute to the layout-grounded T2I branch. In addition, we use virtual try-on data from VITON-HD [choi2021viton] and DressCode-MR[chong2025fastfit] as auxiliary sources, converting them into the same reference-conditioned format to further enrich the subject and appearance diversity of ConsistCompose3M.
Pipeline 1: Layout Diversification via Controlled Outpainting. As shown in Pipeline 1 of Fig. S7, this pipeline broadens the spatial distribution of bounding-box positions and scales in the single-reference setting using image pairs from Subjects200K [tan2025ominicontrol], obtained through OmniControl’s side-by-side paired generation. For each sample, we first generate a dense caption $p^{\text{raw}}_{\text{dense}}$ for the reference image using GPT-4o [hurst2024gpt]. The reference image $R_{i}$ is then resized by a sampled side-length factor $r\sim\mathcal{U}(0.6,0.8)$ and placed onto a blank canvas at a random valid location, thereby defining a normalized bounding box $b\in[0,1]^{4}$.
@@
Finally, the three pipelines collectively yield, for each constructed sample, a scene-level dense description and the corresponding subject–bounding-box associations. We then convert the raw dense sentence $p^{\text{raw}}_{\text{dense}}$ into a layout-grounded, multi-reference prompt $p_{\text{LELG}}$ by inserting explicit bounding-box and source-image tags. For each matched subject, we take its normalized bounding box $x_{1},y_{1},x_{2},y_{2}]$ together with the index of its reference image, and employ GPT-4o [hurst2024gpt] with the instruction template in Fig. S12 to insert a tag of the form $x_{1},y_{1},x_{2},y_{2}]</bbox$> from image $k$ immediately after the first noun-phrase mention of that subject in $p^{\text{raw}}_{\text{dense}}$. By integrating reference-guided synthesis, detection-and-matching–based filtering, and LLM-driven tag insertion, these pipelines produce layout-grounded samples exhibiting rich nested and attributive configurations.
@@ Appendix D Supplementary Results and Detailed Analysis
General Understanding Capabilities. An important consideration is whether layout-aware training compromises general multimodal understanding, which we investigate through additional quantitative ablation studies across two benchmark suites. For fine-grained generation control (Table S7), our Alignment and SFT (incorporating LLM-based rewriting) attain an overall score of 0.88—ranking first among compared methods—with state-of-the-art position control performance (0.84) and strong attribute consistency (0.79–0.80). This confirms that precise layout adherence does not sacrifice control granularity. For general understanding benchmarks (Table S8), our Alignment variants (with/without coordinate data) achieve  81.4 on MMBench, comparable to InternVL3-8B (81.7) [Zhu2025InternVL3EA] and Bagel (81.4), validating no degradation in basic visual understanding capabilities. Notably, on MMMU (assessing complex reasoning), Alignment with coordinate data improves from 39.4 (without coordinate data) to 42.3. This suggests that coordinate data may enhance spatial reasoning, representing an advantage over specialized layout models that often trade generalizability for layout control.
@@
Extended DreamBench Results. To assess generalization beyond layout control, we additionally report DreamBench results (Tab. S9, Figs. S15 and S16). Alignment (w/ Coord Data) outperforms its non-coordinate counterpart (0.611 vs. 0.565 in the DreamBench single-subject DINO identity metric), and SFT (w/ Coord Data) further reaches 0.677 for single-subject and 0.506 for multi-subject generation, surpassing baselines such as OmniGen2 [wu2025omnigen2explorationadvancedmultimodal] and narrowing the gap to top-performing models (Flux Kontext dev [labs2025flux], GPT-4o [hurst2024gpt]). Qualitative examples in Figs. S15 and S16 corroborate these gains, showing strong identity preservation and prompt adherence for single subjects, as well as consistent identities with natural interactions in multi-subject scenes. Together, these results indicate that our method attains strong layout control without compromising general understanding or broad generative capability.